OpenStudy (anonymous):

Find the polynomial function with roots 1, -2, and 5.

jhonyy9 (jhonyy9):

so if you subtract coclusion from the quadratic equation that knowing these roots can be rewriting ,finding from (x-x_1)(x-x_2) than x_1 and x_2 are these roots - so than for a polynomial function ,when we know these 3 roots of this function ,i think that will can find it from (x-1)(x+2)(x-5)=(x^2 +x -2)(x-5)=x^3 -5x^2 +x^2 -5x -2x +10 = =x^3 -4x^2 -7x +10